Factor the polynomial completely
6x^3 + 30x^2y^2+36xy^4

Answer by pvrbrahmam(12)

Q:Factorize 6x^3 + 30x^2y^2+36xy^4
Sol: 6x^3 + 30x^2y^2+36xy^4 = 6x(x^2 +5xy^2 +6y^4)= 6x(x^2 +3xy^2+2xy^2+6y^4)
= 6x(x^2 +3xy^2+2xy^2+6y^4)= 6x(x(x + 3y^2) + 2y^2(x +3y^2))
= 6x((x+ 2y^2)(x + 3y^2))